# Analytical results for the Quantum Alternating Operator Ansatz with Grover Mixer

### v2

@@ VI Conclusions
In the present work, we apply the statistical approach of Headley and Wilhelm [27] on GM-Th-QAOA, obtaining an expression for the expectation value with complexity independent of the number of layers. With the expression, we first solve the conjecture of the threshold curve and then get bounds of different natures, including on the statistical quantities of quantile and the standard score, and on the minimum number of layers required to get a fixed approximation ratio. The bound on the quantile is of particular interest since it reflects explicitly a quadratic Grover-style speed-up. Subsequently, we generalize the GM-Th-QAOA bounds to the general Grover-based QAOA framework, achieving the same asymptotic performance and obtaining the main contribution of this work, that is, the formal establishment that the Grover mixer has the performance bounded by the quadratic bound of the unstructured search problem. That limiting can be severe for combinatorial optimization, as evidenced by the application of Max-Cut on the complete bipartite graph, which requires an exponential number of layers to maintain constant performance. In this way, to get significant results with QAOA, especially in the NISQ era, it is essential for the algorithm to explore the structure of the optimization problems. Indeed, recall the numerical evidence of Golden et al. [14] that suggests the possibility of exponential gain of QAOA with structure-dependent mixer over Grover mixer variants. Thus, research should be directed toward understanding the mechanisms by which different types of mixers can benefit from the structure of particular problems, a path opened by Headley [34] with the statistical approach on the transverse field mixer and the line mixer.
@@
Yet in the Grover mixer context, there are still open questions. Firstly, it would be interesting to decide whether GM-Th-QAOA is the best Grover-based QAOA for all possible instances, or at least whether GM-Th-QAOA outperforms GM-QAOA always, confirming the numerical evidence. Intuitively, it is reasonable to think that the most efficient agnostic-structure method possible is to compile the cost function on a binary function and perform Grover’s algorithm. The results and insights of the present work indicate that this can be the case. However, formal proof is still needed. Moreover, one can ask whether GM-QAOA even reflected the quadratic Grover speed-up in the sense of Theorem 3. However, insights would be needed to answer that question analytically since direct analytical treatment to bound Eq. (30) is infeasible. Finally, the application of the maximum amplification bound to more graph classes on Max-Cut and other combinatorial optimization problems would be welcome. This way, we could know how common is the need for exponential growth on the layers. Fortunately, explicit knowledge of the distribution is not necessarily required to find the asymptotic behavior and establish how the performance scales, as was the case of complete bipartite graphs on Max-Cut.
@@ Acknowledgements.
G.A.B. thanks the financial support of CAPES/Brazil. F.L.M. thanks the financial support of CNPq/Brazil, grant number 407296/2021-2, and Latvian Quantum Initiative. The authors thank R. Portugal and F. Santos for helpful discussions.
@@ Appendix A Proof of Theorem 2
where $B_{2n}$ is the Bernoulli number. Note that both expansions are non-zero for terms with the same order, which allows a direct comparison. We are interested in establishing that all terms of $\tan^{2}{(x)}+2\sin^{2}{(x)}$ expansion are non-negative. The sine squared expansion alternates the sign, having a negative sign for $n$ even, while tangent squared expansion has all positive terms. Then, is enough to show that all even $n$ terms of $\tan^{2}{(x)}$ expansion are equal or greater than the absolute value of the respective terms of $2\sin^{2}{(x)}$. Comparing both expansions on Eq. (62), for $n=2$, they are the same and followed by an induction argument that our claim holds for $n>2$.

### v3

@@ VI Conclusions
In the present work, we apply the statistical approach of Headley and Wilhelm [28] on GM-Th-QAOA, obtaining an expression for the expectation value with complexity independent of the number of layers. With the expression, we first solve the conjecture of the threshold curve and then get bounds of different natures, including on the statistical quantities of quantile and the standard score, and on the minimum number of layers required for the algorithm to guarantee some fixed approximation ratio. The bound on the quantile is of particular interest since it reflects explicitly a quadratic Grover-style speed-up. Subsequently, we generalize the GM-Th-QAOA bounds to the general Grover-based QAOA framework. We derive these findings from the result—established through a proof by contradiction with the optimality of the unstructured search problem—that the probability of measuring states on Grover-based QAOA is bounded by a quadratic growth on the number of layers. The findings showed that Grover-based QAOA achieves, at most, the same asymptotic performance as GM-Th-QAOA. Consequently, we obtain the main contribution of this work: the formal establishment that the Grover mixer’s performance is bounded by the quadratic bound of the unstructured search problem, confirming previous conjectures in the literature [13, 14]. That limiting can be severe for combinatorial optimization, as evidenced by the application of Max-Cut on the complete bipartite graph, which we conclude by an argument using the asymptotic decay rate of the probability distribution that it requires an exponential number of layers to maintain constant performance and, therefore, we cannot have a polynomial-time algorithm that guarantees a fixed approximation ratio. This highlights that in order to get significant results with QAOA, especially in the NISQ era, it is essential for the algorithm to explore the structure of the optimization problems. Indeed, recall the numerical evidence of Golden et al. [14] that suggests the possibility of exponential gain of QAOA with structure-dependent mixer over Grover mixer variants. Thus, research should be directed toward understanding the mechanisms by which different types of mixers can benefit from the structure of particular problems, a path opened by Headley [35] with the statistical approach on the transverse field mixer and the line mixer.
For a more accurate picture of the limitation of QAOA with Grover mixer, future works would consider the application of the maximum amplification bound to more graph classes on Max-Cut and other combinatorial optimization problems, and so estimate how common the need for exponential growth on the layers to achieves a target approximation ratio. Fortunately, explicit knowledge of the distribution is not necessarily required to find the asymptotic behavior and establish how the performance scales—as was the case of complete bipartite graphs on Max-Cut—which could make the task much easier. A case of particular interest is the investigation of classes of instances where QAOA with another mixer besides Grover mixer has some known guarantee on the approximation ratio, such as Max-Cut on $3$-regular graphs [4, 30] and Sherrington-Kirkpatrick model at infinity size [32], both for traverse field mixer. If the maximum amplification bound requires exponential resources to maintain constant performance for a case of this type, the direct comparison would underscore the limitation of the Grover mixer as a mixer for QAOA.
@@
Yet in the Grover mixer context, there are at least two open questions. Firstly, it would be interesting to decide whether GM-Th-QAOA is the best Grover-based QAOA for all possible instances, or at least whether GM-Th-QAOA outperforms GM-QAOA always, confirming the numerical evidence. Intuitively, it is reasonable to think that the most efficient agnostic-structure method possible is to compile the cost function on a binary function and perform Grover’s algorithm. The results and insights of the present work indicate that this can be the case. However, formal proof is still needed. Secondly, one can ask whether GM-QAOA even reflected the quadratic Grover-style speed-up in the sense of Theorem 3. However, insights would be needed to answer that question analytically since direct analytical treatment to bound Eq. (34) is infeasible.
@@ Acknowledgements.
G.A.B. thanks the financial support of CAPES/Brazil. F.L.M. thanks the financial support of CNPq/Brazil grant 407296/2021-2, and Latvian Quantum Initiative project 2.3.1.1.i.0/1/22/I/CFLA/001. The authors thank R. Portugal and F. Santos for helpful discussions.
@@ Appendix A Proof of Theorem 2
where $B_{2n}$ is the Bernoulli number. Note that both expansions are non-zero for terms with the same order, which allows a direct comparison. We are interested in establishing that all terms of $\tan^{2}{(x)}+2\sin^{2}{(x)}$ expansion are non-negative. The sine squared expansion alternates the sign, having a negative sign for $n$ even, while tangent squared expansion has all positive terms. Then, it is enough to show that all even $n$ terms of $\tan^{2}{(x)}$ expansion are equal or greater than the absolute value of the respective terms of $2\sin^{2}{(x)}$. Comparing both expansions on Eq. (62), for $n=2$, they are the same and followed by an induction argument that our claim holds for $n>2$.
@@ Appendix B Proof of Lemma 1
$\operatorname{E}[Y_{\leq T}]=\frac{G_{Y}(T)}{F_{Y}(T)},\ \operatorname{E}[Y_{\leq T}^{2}]=\frac{\sigma^{2}_{\leq T}}{F_{Y}(T)},\ \operatorname{E}[Y_{>T}]=-\frac{G_{Y}(T)}{1-F_{Y}(T)},\ \operatorname{E}[Y_{>T}^{2}]=\frac{\sigma^{2}_{>T}}{1-F_{Y}(T)},$ (64)
Combining both bounds gives
$\frac{|G_{Y}(T)|}{\sigma}\leq\sqrt{F_{Y}(T)(1-F_{Y}(T))},$ (66)
and the equality is hit on the limit of Jensen’s inequality if and only if both $Y_{\leq T}$ and $Y_{>T}$ are degenerate distributions.
Appendix C Proof of Lemma 2
@@
Xie et al.  [55], after the first preprint version of our paper, introduced an alternative proof for Lemma 2 using mathematical optimization techniques to identify critical points through partial derivatives.
Appendix D Proof of Theorem 6
with $F_{X_{1}}(\tau_{2})=1/(2r+1)^{2}$. Now, we can apply the same proceeding of Lemma 1 on the random variables $X_{\leq\tau_{2}}$, the random variable $X_{1}$ given $X_{1}\leq\tau_{2}$; and $X_{\geq\tau_{2}}$, the random variable $X_{2}$ given $X_{2}\geq\tau_{2}$, to get the same conclusion. Then, take binary function on Eq. (77) gives $C(r)\leq 2\sqrt{r(r+1)}$.